# Johannes Matzat        Aiko Schmeißer

### v3

@@ 4 Empirical Strategy
Alternative source of variation. Lastly, we describe our DiD-IV approach, which complements the DiD strategy with arguably exogenous variation in union support driven by spikes in work-related fatalities. After the NLRB accepts a petition to hold a union election, it sets the timeline of the unionization process and fixes an election date. Any random unexpected shocks between petition and election that shift union support are then potential candidates for an instrument. We focus on sector-level fatal work accidents in the 30 days before a union election. [^27] The median time between petition and election in our sample is 45 days. Less than 5% of all elections are held within 30 days after the petition. Safety at work is a fundamental concern to all workers, especially when one’s life is in danger. Work-related fatalities are unfortunately still common in the United States. In 2019, the Occupational Safety and Health Administration (OSHA) reported 1,943 deaths at work, more than 5 per day on average. Unions often campaign on safety issues and are found to improve safety conditions at the workplace (e.g., AFL-CIO, 2022; Hagedorn et al., 2016; Li et al., 2022).
@@ 5.1 Main Results
Figure 2 reports our main results for the effect of unionization on employees’ campaign contributions, using the DiD and event study models (1) and (2). We start with the effects on the (IHS-transformed) total contribution amounts, which are depicted on the left-hand side of the figure. The upper panel plots the results for all employees in an establishment. Note the absence of any significant differential trends between establishments winning and establishments losing the union election in the three cycles (six years) before the election. The effect of unionization on the amount of contributions is small and insignificant in all post-election periods, but we see a moderate spike in contributions in the cycle of the union election (which we are not able to estimate precisely, though). Differentiating between contributions made by workers and managers in the lower panels highlights that workers drive the increase in contributions. The event-study results indicate that unionization raises workers’ contributions by 11% in the cycle of the union election (significant at the 5% level). This pattern is consistent with a short-term political mobilization of workers through a successful union campaign at the workplace. Overall, however, the DiD coefficients indicate that there is no significant average effect on the amount of contributions over the three cycles after a union election.
@@ 6.2 Ideological versus Strategic Motives
Differentiating candidates by union support. We continue by examining whether the partisan shift in employee contributions can be solely explained by a shift from candidates not supported to candidates supported by the union organization. To identify union-supported candidates, we have matched donations from PACs associated with the union organization that was on the ballot in a given union election (including PACs of local union branches). Appendix Table A.11 reports for each union organization in our sample the share of contributions to Democratic (as opposed to Republican) candidates. On average, union PACs give 94% of their donations to Democrats, which demonstrates the strong link between labor unions and the Democratic Party. In Table 6, we use this information to estimate the effects of unionization on employees’ contribution amounts to candidates supported and not supported by the respective union in a cycle. Unions may encourage their members to contribute to the same candidates as their PAC, either directly through providing information about these candidates or indirectly through employees contributing to the union PAC and observing which candidates are supported by the union PAC. While the estimated coefficients for the effect of unionization on employees’ contribution amounts to union-supported candidates are positive, they are small and insignificant. [^40] When taking the difference in the amounts going to union-supported and non-union-supported candidates, the effect of unionization is significant only for contributions from managers and only at the 10% level (not shown). Moreover, we find a shift in contributions from Republican to Democratic candidates that remains significant even after conditioning on whether or not candidates are supported by the union. In other words, among both union-supported and non-union-supported candidates we observe significant effects on contribution amounts to Democratic relative to Republican candidates. These results suggest a limited role for unions’ direct influence in channeling employees’ contributions toward candidates in which the union has strategic interests. Instead, unionization appears to cause a more general leftward shift in employees’ donation behavior.
@@ B.1 Union Election Data
  • For multiple entries that reflect elections where more than one union were on the ballot or where different worker groups formed different bargaining units, we follow Frandsen (2021) and retain only the entry with the largest union vote share.

### v4

@@ 4 Empirical Strategy
Alternative source of variation. Lastly, we describe our DiD-IV approach, which complements the DiD strategy with arguably exogenous variation in union support driven by spikes in work-related fatalities. After the NLRB accepts a petition to hold a union election, it sets the timeline of the unionization process and fixes an election date. Any random unexpected shocks between petition and election that shift union support are then potential candidates for an instrument. We focus on sector-level fatal work accidents in the 30 days before a union election. [^27] The median time between petition and election in our sample is 45 days. Less than 5% of all elections are held within 30 days after the petition. Safety at work is a fundamental concern to all workers, especially when one’s life is in danger. Work-related fatalities are unfortunately still common in the United States. In 2019, the Occupational Safety and Health Administration (OSHA) reported 1,943 deaths at work, more than 5 per day on average. Unions often campaign on safety issues and are found to improve safety conditions in the workplace (e.g., AFL-CIO, 2022; Hagedorn et al., 2016; Li et al., 2022).
@@ 5.1 Main Results
Figure 2 reports our main results for the effect of unionization on employees’ campaign contributions, using the DiD and event study models (1) and (2). We start with the effects on the (IHS-transformed) total contribution amounts, which are depicted on the left-hand side of the figure. The upper panel plots the results for all employees in an establishment. Note the absence of any significant differential trends between establishments winning and establishments losing the union election in the three cycles (six years) before the election. The effect of unionization on the amount of contributions is small and insignificant in all post-election periods, but we see a moderate spike in contributions in the cycle of the union election (which we are not able to estimate precisely, though). Differentiating between contributions made by workers and managers in the lower panels highlights that workers drive the increase in contributions. The event-study results indicate that unionization raises workers’ contributions by 11% in the cycle of the union election (significant at the 5% level). This pattern is consistent with a short-term political mobilization of workers through a successful union campaign in the workplace. Overall, however, the DiD coefficients indicate that there is no significant average effect on the amount of contributions over the three cycles after a union election.
@@ 6.2 Ideological versus Strategic Motives
Differentiating candidates by union support. We continue by examining whether the partisan shift in employee contributions can be solely explained by a shift from candidates not supported to candidates supported by the union organization. To identify union-supported candidates, we have matched donations from PACs associated with the union organization that was on the ballot in a given union election (including PACs of local union branches). Appendix Table A.11 reports for each union organization in our sample the share of contributions to Democratic (as opposed to Republican) candidates. On average, union PACs give 94% of their donations to Democrats, which demonstrates the strong alliance between labor unions and the Democratic Party (Dark, 2001). In Table 6, we use this information to estimate the effects of unionization on employees’ contribution amounts to candidates supported and not supported by the respective union in a cycle. Unions may encourage their members to contribute to the same candidates as their PAC, either directly through providing information about these candidates or indirectly through employees contributing to the union PAC and observing which candidates are supported by the union PAC. While the estimated coefficients for the effect of unionization on employees’ contribution amounts to union-supported candidates are positive, they are small and insignificant. [^40] When taking the difference in the amounts going to union-supported and non-union-supported candidates, the effect of unionization is significant only for contributions from managers and only at the 10% level (not shown). Moreover, we find a shift in contributions from Republican to Democratic candidates that remains significant even after conditioning on whether or not candidates are supported by the union. In other words, among both union-supported and non-union-supported candidates we observe significant effects on contribution amounts to Democratic relative to Republican candidates. These results suggest a limited role for unions’ direct influence in channeling employees’ contributions toward candidates in which the union has strategic interests. Instead, unionization appears to cause a more general leftward shift in employees’ donation behavior.
@@ B.1 Union Election Data
  • For multiple entries that reflect elections where more than one union was on the ballot or where different worker groups formed different bargaining units, we follow Frandsen (2021) and retain only the entry with the largest union vote share.